OpenStudy (anonymous):

Is the following true of false? d^2/dx^2=(x^2(cosx))= 2cos(x)-4xsin(x)-x^2cos(x)

OpenStudy (loser66):

post the original problem, please. What is f(x)??

OpenStudy (anonymous):

This is the original question

OpenStudy (freckles):

oh I think she didn't mean to put an equal sign between the operator and what she was trying to differentiate

OpenStudy (anonymous):

thx freckle for helping me

OpenStudy (anonymous):

there is no f(\(x\))

OpenStudy (freckles):

\[\frac{d^2}{dx^2}(x^2 \cos(x))\] I think this is what she meant by the original problem

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Yes thats right

OpenStudy (freckles):

we first need to find the first derivative and you used the product rule? can I see what you got for this step \[\frac{d}{dx}(\frac{d}{dx}(x^2\cos(x)))\]

OpenStudy (anonymous):

No its asking if what they got for the problem after applying chain rule is true or false.

OpenStudy (freckles):

ok well we can differentiate twice to see

OpenStudy (anonymous):

And I think it is true because when you apply chain rule you do get what they got

OpenStudy (loser66):

yup

OpenStudy (anonymous):

Okay thank you!